find the missing values
it a cone witha voulme= 92.316
radius 42
i need to find the height

Answer by stanbon(75887)

find the missing values
it a cone witha voulme= 92.316
radius 42
i need to find the height
--------------------------------
V = (1/3)Bh where B = pi*r^2
---
V = (1/3)(pi)r^2h
92.316 = (1/3)(pi)(42^2)h
(3*92.316)/(pi*42^2) = h
h = 0.04997 units
===========================
Cheers,
Stan H.